One number is 800 less than three times the other, and their sum is 1200. Find the numbers.
Please show me the solutions. As simplify as it can be...

Answer by mananth(16949)

let one number be x
the other number will be 1200-x. ( Because the sum of both is 1200)
..
one number is 800 less than the 3 times the other.
The other number is 1200-x
3 times other = 3(1200-x)
less 800
= 3(1200-x) -800
x=3(1200-x)-800
x=3600-3x-800
x+3x= 2400
4x = 2400
divide by 4
x= 2400/4
x=600
if one number is 600, then other is 1200-x=600.